# Video LLMs for Temporal Reasoning in Long Videos

### v1

@@ 5 Experiments
Tasks and Datasets. Our TemporalVLM model is evaluated on a few temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. For fair comparisons with TimeChat [40], we fine-tune our model on a subset [^1] We could not download the YT-Temporal [60] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [40] and Valley [32] datasets. The compiled dataset contains a total of 142K long videos with diverse activities. We follow TimeChat [40] to use 6 instructions (generated by GPT-4 [1]) per task and convert the video annotations into chat samples. In the zero-shot setting, we evaluate on YouCook2 [63] for dense video captioning, Charades-STA [10] for temporal video grounding, QVHighlights [25] for video highlight detection, and our IndustryASM dataset for temporal action segmentation.
Implementation Details. We follow TimeChat [40] to use LLaMA-2 7B [47] as the LLM, and ViT-G/14 from EVA-CLIP [44] as the image encoder. We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [61] checkpoints to initialize the video Q-Former. The parameters of the BiLSTM module and projection layers are initialized randomly. The LLM is fine-tuned using LoRA [16] with rank 32. We use 6 clips and sample 96 frames from each clip. We fine-tune our model on a subset of TimeIT [40] and Valley [32] for 7 epochs on a single 8-A100 (80 GB) machine. The parameters of the LLM and image encoder remain frozen throughout the fine-tuning. Due to space limits, please refer to the supplementary documentation for more details.
Competing Methods. We compare our model against state-of-the-art video LLMs, including VideoChat-Embed [27], Video-LLaMA [61], Valley [32], TimeChat [40], and LongVLM [55]. Since the fine-tuned LongVLM model is not available, we replace our time-aware clip encoder and BiLSTM modules with their local and global feature extraction modules respectively while keeping the same image encoder and LLM for fair comparisons, yielding our reimplemented LongVLM model which we use for evaluation.
@@
  • Temporal video grounding: To gauge the effectiveness in temporally grounding user queries, we use R@1(IoU = $\alpha$), which tells the percentage of predictions with an IoU greater than $\alpha$ compared to the ground truth.
@@
  • Temporal action segmentation: We report the framewise accuracy (Acc), F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$ (F1@$\{10,25,50\}$), and Edit distance (Edit). We use the ratio of intersection over union (IoU) to determine the overlapping thresholds.
@@ 5.1 Zero-Shot Results
We first evaluate the performance of all methods in the zero-shot setting. Tab. 1 presents the zero-shot results.
@@ 5.1.3 Temporal Video Grounding Results
It is evident from Tab. 1 that our TemporalVLM model achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7). Specifically, it outperforms TimeChat [40] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [55] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively. The above results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our time-aware clip encoder extracts fine-grained cues for temporal reasoning within each clip, while our BiLSTM module captures long-range temporal relationships between multiple clips to detect the given event within the video effectively. The lack of a time-aware encoder significantly reduces the performance of LongVLM [55], while TimeChat [40] suffers from a lack of fine-grained information.
@@ 5.2 Supervised Results
We now evaluate the performance of all methods in the supervised setting.
@@ 5.2.2 Temporal Action Segmentation Results
Model F1@10 F1@25 F1@50 Edit Acc
TimeChat [40] 11.1 8.5 4.4 32.7 46.0
LongVLM [55] 17.1 13.3 7.3 47.3 47.6
TemporalVLM (Ours) 22.3 18.3 11.1 56.2 52.9
@@
We further evaluate all methods on our IndustryASM dataset for the temporal action segmentation task. Since the activities in IndustryASM are significantly different from those in TimeIT [40] and Valley [32], we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. It is evident from Tab. 3 that our TemporalVLM model achieves the best results across all metrics. For example, it outperforms TimeChat [40] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [55] by +8.9 and +5.3 on Edit and Acc respectively. The above results confirm the advantages of our time-aware and multi-level features for temporal action segmentation. In addition, TimeChat [40] obtains the worst performance, which is likely because temporal action segmentation requires fine-grained information to accurately predict the boundaries of each action and both TemporalVLM and LongVLM [55] utilize multi-clip feature extraction whereas TimeChat [40] treats the entire video as a single clip.
@@ 5.3 Ablation Results
Model SODA_c CIDEr F1 Score
Average Pooling 1.7 6.0 16.1
LSTM 2.8 9.4 18.2
BiLSTM 3.4 13.2 20.0
Number of Clips SODA_c CIDEr F1 Score
@@
Below we study the effects of the BiLSTM module and the number of short-term clips. For this purpose, we fine-tune multiple configurations of our TemporalVLM model on Youcook2 [63].
5.3.1 Impacts of the BiLSTM Module
In Tab. 4, we conduct an ablation study by replacing our BiLSTM module, which learns temporal relationships in both directions, with either a standard LSTM, which extracts temporal dependencies in the forward pass only, or average pooling operation, which is not trainable and employed in prior works, e.g., [55], for aggregating global features. From the results, it is clear that using BiLSTM for global feature aggregation yields the best results, which confirms the advantages of using BiLSTM over using LSTM or average pooling in our model.
@@ 5.4 Qualitative Results
In the following, we discuss some qualitative results of TemporalVLM, TimeChat [40], and LongVLM [55].
@@ 6 Conclusion
We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos. Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features. As a result, the extracted features not only are time-sensitive but also contain both local and global cues. In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes. Our dataset is recorded on factory floors and includes action and timestamp labels provided by industrial engineers for time and motion studies and temporal action segmentation evaluation. Lastly, we conduct extensive experiments on datasets of long videos, i.e., TimeIT [40] and IndustryASM, to show our superior results over previous methods across temporal reasoning and fine-grained understanding tasks such as dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
@@ Appendix A Supplementary Material
In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs. A.1 and A.2 respectively. Next, Sec. A.3 compares the sizes of our model and TimeChat [40] in terms of the number of learnable parameters. We then present several qualitative results in Sec. A.4, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results. Finally, we discuss the limitations, future works, and societal impacts of our work in Sec. A.5.
@@ A.1.1 Dataset Statistics
A.1.2 Chat Samples
Our IndustryASM videos are annotated with action names and action timestamps by industrial engineers. We manually write the user instructions for the temporal action segmentation task and generate the ground truth responses by using the action names and action timestamps. Fig. 6 illustrates examples of our instruction and ground truth response for the temporal action segmentation on an IndustryASM video.
@@ A.2 Implementation Details
We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab. 7. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. The output of the clip encoder contains a time-aware representation of each clip. We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM. To the best of our knowledge, our work is the first to incorporate LSTM into video LLMs.
@@ A.3 Comparison on Model Sizes
A.4 Qualitative Results
@@
A.5 Discussions
A.5.1 Limitations and Future Works
Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex/dense temporal reasoning tasks and videos with extreme durations. Our future works will enhance TemporalVLM for tackling complex/dense temporal reasoning tasks and videos with extreme durations. In addition, we will explore video LLMs with both spatial [4, 5] and temporal reasoning capabilities.
A.5.2 Societal Impacts

### v2

@@ 5 Experiments
Tasks and Datasets. Our TemporalVLM model is evaluated on a few temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. For fair comparisons with TimeChat [35], we fine-tune our model on a subset [^1] We could not download the YT-Temporal [54] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [35] and Valley [27] datasets. The combined dataset contains a total of 142K long videos with diverse activities. We follow TimeChat [35] to use 6 instructions (generated by GPT-4 [1]) per task and convert the video annotations into chat samples. We evaluate on YouCook2 [58] for dense video captioning, Charades-STA [8] for temporal video grounding, QVHighlights [19] for video highlight detection, and our IndustryASM dataset for temporal action segmentation.
Implementation Details. We follow TimeChat [35] to use LLaMA-2 7B [42] as the LLM, and ViT-G/14 from EVA-CLIP [39] as the image encoder. We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [56] checkpoints to initialize the video Q-Former. The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer. The parameters of the BiLSTM module and projection layers are initialized randomly. The LLM is fine-tuned using LoRA [12] with rank 32. We use 6 clips and sample 96 frames from each clip. We fine-tune our model on a subset of TimeIT [35] and Valley [27] for 7 epochs on a single 8-A100 (80 GB) machine. The parameters of the LLM and image encoder remain frozen throughout the fine-tuning. Please see our supplementary material for more details.
Competing Methods. We compare our model against state-of-the-art video LLMs, including VideoChat-Embed [21], Video-LLaMA [56], Valley [27], TimeChat [35], and LongVLM [49]. Since LongVLM [49] is not trained on TimeIT [35] and Valley [27] and it is a multi-model pipeline (TimeChat [35] and TemporalVLM are end-to-end models), we replace our time-aware clip encoder and BiLSTM modules with their local and global feature extraction modules respectively while keeping the same image encoder and LLM, yielding our reimplemented LongVLM model which we use for evaluation. Nevertheless, we compare with the original LongVLM model [49] on general video understanding in our supplementary material.
@@
  • Temporal video grounding: To gauge the effectiveness of temporal video grounding, we use R@1(IoU = $\alpha$), which is the percentage of predictions with an intersection over union greater than $\alpha$ compared to the ground truth.
@@
  • Temporal action segmentation: We report the framewise accuracy, F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$, and Edit distance. We use the intersection over union ratio to find the overlapping thresholds.
@@ 5.1 Zero-Shot Results
We first evaluate all methods in the zero-shot setting and present the results in Tab. 1.
@@ 5.1.3 Temporal Video Grounding Results
It is evident from Tab. 1 that TemporalVLM achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7). Specifically, it outperforms TimeChat [35] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [49] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively. The results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our time-aware clip encoder extracts fine-grained cues for temporal reasoning within each clip, while our BiLSTM module captures long-range temporal relationships between multiple clips to detect the given event within the video effectively. The lack of a time-aware encoder reduces the performance of LongVLM [49], while TimeChat [35] suffers from a lack of fine-grained information.
@@ 5.2 Supervised Results
We now evaluate all methods in the supervised setting.
@@ 5.2.2 Temporal Action Segmentation Results
Model F1@{10,25,50} Edit Acc
TimeChat [35] {11.1, 8.5, 4.4} 32.7 46.0
LongVLM [49] {17.1, 13.3, 7.3} 47.3 47.6
TemporalVLM (Ours) {22.3, 18.3, 11.1} 56.2 52.9
@@
We further test all methods on our IndustryASM dataset for temporal action segmentation. Since the activities in IndustryASM significantly differ from those in TimeIT [35] and Valley [27], we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. It is evident from Tab. 3 that TemporalVLM performs the best across all metrics. For example, it outperforms TimeChat [35] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [49] by +8.9 and +5.3 on Edit and Acc respectively. The results confirm the benefits of our time-aware and multi-level features for temporal action segmentation.
Multi-Clip Encoder vs. Time-Aware Encoder. Multi-clip encoders in LongVLM [49] and TemporalVLM are crucial to capturing fine-grained cues, leading to superior results over TimeChat [35] on temporal action segmentation in Tab. 3. In contrast, time-aware encoders in TimeChat [35] and TemporalVLM are important to event localization and temporal modeling, resulting in better performance than LongVLM [49] on dense video captioning, video highlight detection, and temporal video grounding in Tabs. 1 and 2. Our TemporalVLM encoder is both multi-clip and time-aware, yielding the best performance across all tasks.
@@ 5.3 Ablation Results
# Clips SODA_c CIDEr F1
@@
Model # Layers SODA_c CIDEr F1
Average Pooling 1 1.7 6.0 16.1
Linear Layer 1 2.4 9.5 17.8
Transformer 2 2.3 6.8 15.6
LSTM 1 2.8 9.4 18.2
BiLSTM 2 3.4 13.2 20.0
Below we study the effects of the number of short-term clips and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [58].
@@ 5.3.1 Impacts of the Number of Short-Term Clips
5.3.2 Impacts of the BiLSTM Module
To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab. 5. From Tab. 5, BiLSTM outperforms average pooling, which is not trained and employed in prior works [28, 49], and traditional LSTM, which only performs the forward pass and hence relies only on past cues. In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues. Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers [56, 35] alone, leading to worse results. This highlights the importance of BiLSTM even when video Q-Formers [56, 35] are being used. Finally, BiLSTM outperforms transformer [43], which we will discuss below.
(Bi)LSTM vs. Transformer. The permutation-invariant nature of the self-attention mechanism in transformer [43] leads to loss in temporal information of the input sequence. Though using positional encoding alleviates this issue, it is inevitable that some temporal information loss occurs [55]. In contrast, (Bi)LSTM processes data one by one and preserves the order of the sequence, making them better suited for temporal modeling. Similar to [55], we also observe in Tab. 5 that linear layer outperforms transformer [43].
@@ 5.4 Qualitative Results
In this section, we discuss some qualitative results.
@@ 6 Conclusion
We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos. Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features. The extracted features not only are time-sensitive but also contain both local and global cues. In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes. Our dataset is recorded on factory floors and includes action and timestamp labels provided by industrial engineers for time and motion studies and temporal action segmentation evaluation. Lastly, we conduct extensive experiments on TimeIT [35] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to incorporate LSTMs into video LLMs.
@@ Appendix A Supplementary Material
In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs. A.1 and A.2 respectively. Next, Sec. A.3 compares the sizes of our model, TimeChat [35], and LongVLM [49], while Sec. A.4 evaluates our model for general video understanding. We then present several qualitative results in Sec. A.5, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results. Finally, we discuss the limitations and societal impacts of our work in Sec. A.6.
@@ A.1.1 Dataset Statistics
A.1.2 Annotation Process
Our IndustryASM videos are annotated with action names and action timestamps by industrial engineers. To ensure quality, two labelers are assigned for each video, i.e., one labeler provides the labels for the video, while another labeler checks the labels. If there are conflicts, both labelers discuss to fix them. In general, roughly 8% of the videos have conflicts and need fixing, resulting in an agreement rate of around 92% between labelers. We manually write the user instructions for the temporal action segmentation task and generate the ground truth responses by using the action names and action timestamps. Fig. 6 illustrates examples of our instruction and ground truth response for the temporal action segmentation on an IndustryASM video.
@@ A.2 Implementation Details
We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab. 7. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. The output of the clip encoder contains a time-aware representation of each clip. We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$. The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
@@ A.3 Comparison on Model Sizes
A.4 General Video Understanding Results
Method Data CI DO CU TU C
Video-LLaMA [56] 10M 1.96 2.18 2.16 1.82 1.79
Video-ChatGPT [28] 100k 2.50 2.57 2.69 2.16 2.20
Valley [27] 234k 2.43 2.13 2.86 2.04 2.45
BT-Adapter [24] 10M 2.68 2.69 3.27 2.34 2.46
LongVLM [49] 100k 2.76 2.86 3.34 2.39 3.11
TemporalVLM 100k 2.88 2.91 3.45 2.50 3.16
Table 9: General video understanding results. Best results are in bold, while second best ones are underlined.
Following LongVLM [49], we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT [28] in Tab. 9. The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C). It is evident from Tab. 9 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model [49].
A.5 Qualitative Results
@@
A.6 Discussions
A.6.1 Limitations
Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex/dense temporal reasoning tasks and videos with extreme durations. Our future works will enhance TemporalVLM for tackling complex/dense temporal reasoning tasks and videos with extreme durations.
A.6.2 Societal Impacts